# Unsupervised Representation Learning from Sparse Transformation Analysis

### v1

@@ 4.3 Fluid-Dynamic OT for Posterior Flow
The divergence-free vector field is supposed to learn the vorticity which can model periodic transformations such as rotation. For the curl-free counterpart, the irrotational vector field defines the steepest descent direction on the potential landscape. As a beneficial inductive bias for the path, we thus would like the potential flow to follow Optimal Transport (OT). When the vector field $\nabla u$ satisfies certain PDEs, the evolution of the probability density can be seen to minimize the $L_{2}$ Wasserstein distance between the source distribution and the target distribution. Formally, we have:
Theorem 4.1 (Benamou-Brenier Formula [90]).
@@
[…]
@@
Solving the above equations by Karush–Kuhn–Tucker (KKT) conditions gives the optimal solution: the Hamilton-Jacobi (HJ) equation ($\partial_{t}u+\nicefrac{{1}}{{2}}||\nabla u||^{2}{=}0$). To enforce the OT property to the potential flow, we place the following PINN constraint:
$\begin{gathered}{\mathcal{L}}_{HJ}=\frac{1}{T}\sum_{t=1}^{T}\sum_{k}{\bm{g}}_{t}^{k}\big{(}\frac{\partial}{\partial t}u^{k}({\bm{z}},t)+\frac{1}{2}||\nabla_{\bm{z}}u^{k}({\bm{z}},t)||^{2}\big{)}^{2}\end{gathered}$ (17)
@@ 4.4 Brownian Motion for Latent Prior Evolution
In line with [12], as we do not assume any prior knowledge of each transformation, we would like to enforce minimally informative priors. This can be achieved by considering the time evolution as Brownian motion, i.e., random trajectories. To this end, we define the potential function $\psi^{k}({\bm{z}})=-D_{k}\log p({\bm{z}}_{t})$ which advects the density $p({\bm{z}})$ through the induced velocity field $\nabla\psi^{k}({\bm{z}})$. Then according to the continuity equation, the prior evolves as:
@@ 4.5 Latent Posterior Time Evolution
TABLE II: Equivariance error $\mathcal{E}_{k}$ and average log-likelihood $\log p({\bm{x}}_{t})$ on Shapes3D [91].
@@
VAE [89] No () 6924.63$\pm$8.92 7746.37$\pm$8.77 4383.54$\pm$9.26 2609.59$\pm$7.41 -11784.69$\pm$4.87
$\beta$-VAE [18] No () 2243.95$\pm$12.48 2279.23$\pm$13.97 2188.73$\pm$12.61 2037.94$\pm$11.72 -11924.83$\pm$5.64
FactorVAE [37] No () 1985.75$\pm$13.26 1876.41$\pm$11.93 1902.83$\pm$12.27 1657.32$\pm$11.05 -11802.17$\pm$5.69
SlowVAE [5] Weak () 1247.36$\pm$12.49 1314.86$\pm$11.41 1102.28$\pm$12.17 1058.74$\pm$10.96 -11674.89$\pm$5.74
TVAE [11] Yes () 1225.47$\pm$9.82 1246.32$\pm$9.54 1261.79$\pm$9.86 1142.01$\pm$9.37 -11475.48$\pm$5.18
PoFlow [32] Yes () 885.46$\pm$10.37 916.71$\pm$10.49 912.48$\pm$9.86 924.39$\pm$10.05 -11335.84$\pm$4.95
LatentFlow [12] Yes () 613.29$\pm$8.93 653.45$\pm$9.48 605.79$\pm$8.63 599.71$\pm$9.34 -11215.42$\pm$5.71
LatentFlow [12] Weak () 690.84$\pm$9.57 717.74$\pm$10.65 681.59$\pm$9.02 653.58$\pm$9.57 -11279.61$\pm$5.89
STA No () 1005.23$\pm$11.79 1171.69$\pm$13.64 928.10$\pm$11.58 894.77$\pm$10.94 -11199.93$\pm$5.93
Figure 5: Traversals using each individual learned flow field on Shapes3D [91]. In the bracket, we indicate the transformation which the traversal results look most like. Each latent flow has separate samples per row transforming from left to right.
@@ 5.1.1 Datasets
We evaluate our method on two widely-used benchmarks for standard representation learning, namely MNIST [82] and Shapes3D [91]. For MNIST, The basic ‘pure’ transformations consist of Scaling, Rotation, and Coloring. For Shapes3D, we use the self-contained four transformation primitives, including Floor Hue, Wall Hue, Object Hue, and Scale. On both datasets, we use our spike and slab prior to generate sequences that are composed of ‘composite’ transformations.
Beyond the toy datasets, we also evaluate our method on challenging Falcol3D and Issac3D [14], two complex large-scale and real-world datasets that contain sequences of different transformations. Specifically, Falcol3D consists of indoor 3D scenes with different lighting conditions and camera positions, while Isaac3D is comprised of various robot-arm movements in dynamic environments. Since the image sequences are short, we do not consider speed variations but only enforce the spike prior to generate data sequences with sparsely-varying transformations.
We further conduct some preliminary experiments of applying our method to real-world video analysis, including autonomous driving videos on Cityscape [16] and behavior videos of social agents on CalMS [15]. Different from the used datasets above, we directly feed raw video sequences as input and let the model discover independent motions.
@@ 5.1.2 Baselines
We compare our method with some representative approaches in the field of disentangled and equivariant representation learning, including LatentFlow [12] and PoFlow [32] which adopt potential flow to evolve the latent samples, Topographic VAE (TVAE) [11] which posses topographic structured latent space, SlowVAE [5] which proposes the sparse Laplacian prior $p({\bm{z}}_{t}|{\bm{z}}_{t-1})=\prod\nicefrac{{\alpha\lambda}}{{2\Gamma(1/\alpha)}}\exp{(-\lambda|z_{t,i}-z_{t-1,i}|^{\alpha})}$, and $\beta$-VAE [18] and FactorVAE [37] which encourage the factorization of the single dimensions of latent samples. We also use the vanilla VAE [89] as a controlled baseline.
@@ 5.1.3 Metrics
We mainly evaluate the baselines using the equivariance error which is defined as $\mathcal{E}_{k}{=}\sum_{t=1}^{T}|{\bm{x}}_{t}-\texttt{Decode}({\bm{z}}_{t})|$ where ${\bm{x}}_{t}$ is the element of sequences of each transformation primitive (e.g., scaling and rotation). Since our method is unsupervised, we inspect the traversal results of each basic vector field $\nabla u^{k}+{\bm{r}}^{k}$ and select the index whose flow looks the most like the target transformation. The average log-likelihood of the sequence is also evaluated on the test set. Besides these two metrics, we also adopt the metric Variational Predictability (VP) score [23] to evaluate the disentanglement performance. Readers are kindly referred to Sec. D.3 of the supplementary for these results.
@@ 5.1.4 Implementation Details
Algorithm 1 presents the training algorithm of our method. In practice, it is hard to learn both components ${{\bm{y}}}_{t}$ and $\tilde{{\bm{g}}}_{t}$ simultaneously from the very beginning as the model could learn to use the speed $\tilde{{\bm{g}}}_{t}$ for choosing vector fields (by tuning the magnitude). To avoid this issue, we divide the training process into two stages. In the first stage, we only train the spike components ${{\bm{y}}}_{t}$ to learn to select the basis vector fields. Then in the second stage, we introduce the slab component $\tilde{{\bm{g}}}_{t}$ into the training to learn the additional control of the transformation speed. We see that this two-stage training strategy can help the optimization of these two components. We leave the rest of the implementation details to Sec. C of the supplementary material.
@@ 5.2.1 Qualitative Results
Fig. 4 displays the traversal results of each learned latent flow under different speeds on MNIST. Our model simultaneously disentangles the transformation categories and speeds into these vector fields in an unsupervised manner. We see that each flow field corresponds to a distinct transformation and further presents a precise control of the transformation speed. When increasing the magnitude of ${\bm{g}}^{k}$, the transformation process will be accelerated, i.e., the object will rotate more degrees, get scaled with a larger factor, and change the hue more. Fig. 5 and 6 present the traversal results and the speed variations on Shapes3D. Our method still allows for disentanglement of the transformation categories and speed. We note that speed control is a major merit of our approach as the explicit control of transformation speed is seldom explored in deep representation learning.
@@ 5.2.2 Quantitative Results
Table I and II present the evaluation results of the equivariance error and log-likelihood on MNIST and Shapes3D, respectively. We see STA achieves very competitive performance against other baselines. Specifically, STA outperforms all the unsupervised approaches by a large margin on equivariance error and rivals PoFlow [32] which requires supervision of each transformation primitive. Moreover, our method yields the highest log-likelihood on the test set, which is likely accounted for by the fact that our method incorporates a sophisticated transformation-centric prior over latent states which matches the statistics of the data. The sparse combination of multiple transformations thus can be seen as a kind of data augmentation. Among the transformations, the rotation has the smallest equivariance error. We expect this gain is largely due to the rotational vector field ${\bm{r}}({\bm{z}})$ introduced by the Helmholtz decomposition.
@@ 5.3.3 Periodic Transformations
Figure 10: Traversals using each individual learned flow field on Falcol3D [14]. In the bracket, we indicate the transformation which the traversal results look most like. Each latent flow has separate samples per row transforming from left to right. The bottom row displays the traversal result generated by the $6$’th latent flow field.
@@
TVAE [11] 11477.81 12568.32 11807.34 11829.33 11539.69 11736.78 11951.45
PoFlow [32] 8312.97 7956.18 8519.39 8871.62 8116.82 8534.91 8994.63
LatentFlow [12] 5798.42 6145.09 6334.87 6782.84 6312.95 6513.68 6614.27
@@ 5.5 Real-world Video Analysis
Figure 11: Traversals using each individual learned flow field on Issac3D [14]. In the bracket, we indicate the transformation which the traversal results look most like. Each latent flow has separate samples per row transforming from left to right.
@@
TVAE [11] 8441.65 8348.23 8495.31 8251.34 8291.70 8741.07 8456.78 8512.09
PoFlow [32] 6572.19 6489.35 6319.82 6188.59 6517.40 6712.06 7056.98 6343.76
LatentFlow [12] 3659.72 3993.33 4170.27 4359.78 4225.34 4019.84 5514.97 3876.01
@@ 5.5.1 Robot Arms and Indoor Scenes
Fig. 10 and 11 show the learned latent flows on Falcol3D and Issac3D [14], respectively. As can be seen from the figures, even on these challenging large-scale datasets, our method still allows for unsupervised disentanglement of complex real-world transformations. Table V and VII compares the equivariance error on the two datasets. Similar to the results on MNIST and Shapes3D, our method achieves very competitive performance against supervised ones. This demonstrates that the proposed sparsity priors also scale up to sequences of complex transformations.
Figure 12: Traversal results of learned latent flows on CalMS [15]. For each latent flow, we display two exemplary sequences, and the flow transforms the image from left to right.
@@ 5.5.2 Agent Behavioral Videos
TABLE VIII: Behavior classification results on CalMS [15].
Method MARS [92] B-Kind [15] Trajectory-LSTM [15] STA
@@
We apply our method to disentangle the complex social interactions of mice on Task $1$ of CalMS [15]. On this dataset, there exist three ground truth interactions, namely ‘investigation’, ‘attack’, and ‘mount’. We thus define $3$ latent flows and let the model learn the interactions from the raw videos. The images are of the resolution $128\times 128$.
@@ 5.5.3 Autonomous Driving Videos
Finally, we take a step further to evaluate our method on Cityscape [16], the challenging real-world autonomous driving videos. We take the sequences of segmentation masks as the training data and downsample the resolution to $64{\times}64$. Fig. 14 displays some exemplary traversals of different latent flows. On this dataset, there are no ground truth generative factors so we may have some reasonable interpretations according to the disentangled transformations: we may interpret latent flow 00 as turning left (the sidewalk region on the right side shrinks), latent flow $1$ as getting closer to the front car (the car region expands), latent flow $2$ as getting away from the front car (the car region shrinks and disappears), and latent flow $3$ as changing the right side from terrain to sidewalk. Notice that this is an initial attempt to apply our method to complex real-world video analysis. Nonetheless, this preliminary experiment demonstrates that our method could have real-world applicability for video understanding.
Figure 14: Traversals results of learned flow fields on downsampled segmentation masks of Cityscape [16]. Each latent flow transforms the image from left to right.

### v2

@@ 4.3 Fluid-Dynamic OT for Posterior Flow
The divergence-free vector field is introduced to model the vorticity component of the transformation dynamics, which is particularly well-suited for capturing periodic transformations such as rotation. These vector fields are solenoidal and naturally describe closed-loop flows, making them a useful inductive bias for encoding cyclic behavior in latent dynamics. In contrast, the curl-free component corresponds to an irrotational vector field, which defines the direction of steepest descent on a scalar potential landscape. As an additional beneficial inductive bias for the path, we thus would like the potential flow to follow Optimal Transport (OT). When the vector field $\nabla u$ satisfies certain PDEs, the evolution of the probability density can be seen to minimize the $L_{2}$ Wasserstein distance between the source distribution and the target distribution. Formally, we have:
Theorem 4.1 (Benamou-Brenier Formula (BBF) [benamou2000computational]).
@@
$\begin{gathered}W_{2}(\mu_{0},\mu_{1})^{2}=\min_{\rho,v}\Big\{\int\int\frac{1}{2}\rho(x,t)|v(x,t)|^{2}\\ \mathop{dx}\mathop{dt}:\frac{\mathop{d}\rho(x,t)}{\mathop{dt}}=-\nabla\cdot(v(x,t)\rho(x,t))\Big\}\end{gathered}$ (15)
@@
[…]
@@
Assumptions and Empirical Observations. One key assumption of the BBF is the regularity of the velocity field. Specifically, $v(x,t)$ is not only square-integrable but also typically assumed to be Lipschitz or Sobolev smooth. Although our model does not explicitly enforce such regularity, the velocity fields are defined as gradients of neural networks. In practice, we use smooth activation functions Tanh and GeLU, which make the resulting vector fields $v(z,t)=\nabla u(z,t)$ continuously differentiable and often empirically Lipschitz over compact regions of the latent space. While this does not constitute a formal guarantee, the observed smoothness provides practical support for the effectiveness of the OT-inspired Hamilton-Jacobi regularization in our framework.
@@ 4.4 Brownian Motion for Latent Prior Evolution
In line with [song2023flow], as we do not assume any prior knowledge of each transformation, we would like to enforce minimally informative priors. This can be achieved by considering the time evolution as Brownian motion, i.e., random trajectories. To this end, we define the potential function $\psi^{k}({\bm{z}})=-D_{k}\log p({\bm{z}}_{t})$ which advects the density $p({\bm{z}})$ through the induced velocity field $\nabla\psi^{k}({\bm{z}})$. Then according to the continuity equation, the prior evolves as:
@@ 4.5 Latent Posterior Time Evolution
TABLE II: Equivariance error $\mathcal{E}_{k}$ and average log-likelihood $\log p({\bm{x}}_{t})$ on Shapes3D [3dshapes18].
@@
[…]
Figure 5: Traversals using each individual learned flow field on Shapes3D [3dshapes18]. In the bracket, we indicate the transformation which the traversal results look most like. Each latent flow has separate samples per row transforming from left to right.
@@ 5.1.1 Datasets
We evaluate our method on two widely-used benchmarks for standard representation learning, namely MNIST [lecun1998mnist] and Shapes3D [3dshapes18]. For MNIST, The basic ‘pure’ transformations consist of Scaling, Rotation, and Coloring. For Shapes3D, we use the self-contained four transformation primitives, including Floor Hue, Wall Hue, Object Hue, and Scale. On both datasets, we use our spike and slab prior to generate sequences that are composed of ‘composite’ transformations.
Beyond the toy datasets, we also evaluate our method on challenging Falcol3D and Issac3D [nie2020semi], two complex large-scale and real-world datasets that contain sequences of different transformations. Specifically, Falcol3D consists of indoor 3D scenes with different lighting conditions and camera positions, while Isaac3D is comprised of various robot-arm movements in dynamic environments. Since the image sequences are short, we do not consider speed variations but only enforce the spike prior to generate data sequences with sparsely-varying transformations.
We further conduct some preliminary experiments of applying our method to real-world video analysis, including autonomous driving videos on Cityscape [cordts2016cityscapes] and behavior videos of social agents on CalMS [sun2021multi]. Different from the used datasets above, we directly feed raw video sequences as input and let the model discover independent motions.
@@ 5.1.2 Baselines
We compare our method with some representative approaches in the field of disentangled and equivariant representation learning, including LatentFlow [song2023flow] and PoFlow [song2023latent] which adopt potential flow to evolve the latent samples, Topographic VAE (TVAE) [keller2021topographic] which posses topographic structured latent space, SlowVAE [klindt2021towards] which proposes the sparse Laplacian prior $p({\bm{z}}_{t}|{\bm{z}}_{t-1})=\prod\nicefrac{{\alpha\lambda}}{{2\Gamma(1/\alpha)}}\exp{(-\lambda|z_{t,i}-z_{t-1,i}|^{\alpha})}$, and $\beta$-VAE [higgins2016beta] and FactorVAE [kim2018disentangling] which encourage the factorization of the single dimensions of latent samples. We also use the vanilla VAE [kingma2013auto] as a controlled baseline.
@@ 5.1.3 Metrics
We mainly evaluate the baselines using the equivariance error which is defined as $\mathcal{E}_{k}{=}\sum_{t=1}^{T}|{\bm{x}}_{t}-\texttt{Decode}({\bm{z}}_{t})|$ where ${\bm{x}}_{t}$ is the element of sequences of each transformation primitive (e.g., scaling and rotation). Since our method is unsupervised, we inspect the traversal results of each basic vector field $\nabla u^{k}+{\bm{r}}^{k}$ and select the index whose flow looks the most like the target transformation. The average log-likelihood of the sequence is also evaluated on the test set. Besides these two metrics, we also adopt the metric Variational Predictability (VP) score [zhu2020learning] to evaluate the disentanglement performance. Readers are kindly referred to Sec. D.3 of the supplementary for these results.
@@ 5.1.4 Implementation Details
Algorithm 1 presents the training algorithm of our method. In practice, it is hard to learn both components ${{\bm{y}}}_{t}$ and $\tilde{{\bm{g}}}_{t}$ simultaneously from the very beginning as the model could learn to use the speed $\tilde{{\bm{g}}}_{t}$ for choosing vector fields (by tuning the magnitude). To avoid this issue, we divide the training process into two stages. In the first stage, we only train the spike components ${{\bm{y}}}_{t}$ to learn to select the basis vector fields. The first stage focuses exclusively on training the spike variables until convergence, which we determine based on two explicit criteria: (1) Sparsity Criterion: The average sparsity ratio (percentage of inactive elements) of the spike variables becomes sufficiently high (i.e, ${\bm{y}}_{t}||_{0}<1+\eta$ where $\eta$ is a small tolerance constant we set to $0.3$). This ensures that the spike variables effectively select the relevant transformation vector fields. (2) Convergence Criterion: The approximation error stabilizes and does not further reduce, confirming that the selected vector fields adequately capture the underlying transformations. After meeting both criteria, we initiate the second training stage and introduce the slab variable $\tilde{{\bm{g}}}_{t}$ into the training to learn the additional control of the transformation speed. We see that this two-stage training strategy can help the optimization of these two components. We leave the rest of the implementation details to Sec. C of the supplementary material.
@@ 5.2.1 Qualitative Results
Figs. 4 displays the traversal results of each learned latent flow under different speeds on MNIST. Our model simultaneously disentangles the transformation categories and speeds into these vector fields in an unsupervised manner. We see that each flow field corresponds to a distinct transformation and further presents a precise control of the transformation speed. When increasing the magnitude of ${\bm{g}}^{k}$, the transformation process will be accelerated, i.e., the object will rotate more degrees, get scaled with a larger factor, and change the hue more. Fig. 5 and 6 present the traversal results and the speed variations on Shapes3D. Our method still allows for disentanglement of the transformation categories and speed. We note that speed control is a major merit of our approach as the explicit control of transformation speed is seldom explored in deep representation learning.
@@ 5.2.2 Quantitative Results
Table I and II present the evaluation results of the equivariance error and log-likelihood on MNIST and Shapes3D, respectively. We see STA achieves very competitive performance against other baselines. Specifically, STA outperforms all the unsupervised approaches by a large margin on equivariance error and rivals PoFlow [song2023latent] which requires supervision of each transformation primitive. Moreover, our method yields the highest log-likelihood on the test set, which is likely accounted for by the fact that our method incorporates a sophisticated transformation-centric prior over latent states which matches the statistics of the data. The sparse combination of multiple transformations thus can be seen as a kind of data augmentation. Among the transformations, the rotation has the smallest equivariance error. We expect this gain is largely due to the rotational vector field ${\bm{r}}({\bm{z}})$ introduced by the Helmholtz decomposition.
@@ 5.3.3 Periodic Transformations
Figure 10: Traversals using each individual learned flow field on Falcol3D [nie2020semi]. In the bracket, we indicate the transformation which the traversal results look most like. Each latent flow has separate samples per row transforming from left to right. The bottom row displays the traversal result generated by the $6$’th latent flow field.
@@
TVAE [keller2021topographic] 11477.81 12568.32 11807.34 11829.33 11539.69 11736.78 11951.45
PoFlow [song2023latent] 8312.97 7956.18 8519.39 8871.62 8116.82 8534.91 8994.63
LatentFlow [song2023flow] 5798.42 6145.09 6334.87 6782.84 6312.95 6513.68 6614.27
@@ 5.5 Real-world Video Analysis
Figure 11: Traversals using each individual learned flow field on Issac3D [nie2020semi]. In the bracket, we indicate the transformation which the traversal results look most like. Each latent flow has separate samples per row transforming from left to right.
@@
TVAE [keller2021topographic] 8441.65 8348.23 8495.31 8251.34 8291.70 8741.07 8456.78 8512.09
PoFlow [song2023latent] 6572.19 6489.35 6319.82 6188.59 6517.40 6712.06 7056.98 6343.76
LatentFlow [song2023flow] 3659.72 3993.33 4170.27 4359.78 4225.34 4019.84 5514.97 3876.01
@@ 5.5.1 Robot Arms and Indoor Scenes
Fig. 10 and 11 show the learned latent flows on Falcol3D and Issac3D [nie2020semi], respectively. As can be seen from the figures, even on these challenging large-scale datasets, our method still allows for unsupervised disentanglement of complex real-world transformations. Table V and VII compares the equivariance error on the two datasets. Similar to the results on MNIST and Shapes3D, our method achieves very competitive performance against supervised ones. This demonstrates that the proposed sparsity priors also scale up to sequences of complex transformations.
Figure 12: Traversal results of learned latent flows on CalMS [sun2021multi]. For each latent flow, we display two exemplary sequences, and the flow transforms the image from left to right.
@@ 5.5.2 Agent Behavioral Videos
TABLE VIII: Behavior classification results on CalMS [sun2021multi].
Method MARS [segalin2021mouse] B-Kind [sun2021multi] Trajectory-LSTM [sun2021multi] STA
@@
Figure 14: Traversals results of learned flow fields on downsampled segmentation masks of Cityscape [cordts2016cityscapes]. Each latent flow transforms the image from left to right.
We apply our method to disentangle the complex social interactions of mice on Task $1$ of CalMS [sun2021multi]. On this dataset, there exist three ground truth interactions, namely ‘investigation’, ‘attack’, and ‘mount’. We thus define $3$ latent flows and let the model learn the interactions from the raw videos. The images are of the resolution $128\times 128$.
@@ 5.5.3 Autonomous Driving Videos
Finally, we take a step further to evaluate our method on Cityscape [cordts2016cityscapes], the challenging real-world autonomous driving videos. We take the sequences of segmentation masks as the training data and downsample the resolution to $64{\times}64$. Fig. 14 displays some exemplary traversals of different latent flows. On this dataset, there are no ground truth generative factors so we may have some reasonable interpretations according to the disentangled transformations: we may interpret latent flow 0 as turning left (the sidewalk region on the right side shrinks), latent flow $1$ as getting closer to the front car (the car region expands), latent flow $2$ as getting away from the front car (the car region shrinks and disappears), and latent flow $3$ as changing the right side from terrain to sidewalk. Notice that this is an initial attempt to apply our method to complex real-world video analysis. Nonetheless, this preliminary experiment demonstrates that our method could have real-world applicability for video understanding.
@@ 6 Conclusion
[Uncaptioned image] Yue Song received the B.Sc. cum laude from KU Leuven, Belgium and the joint M.Sc. summa cum laude from the University of Trento, Italy and KTH Royal Institute of Technology, Sweden, and the Ph.D. summa cum laude from the Multimedia and Human Understanding Group (MHUG) at the University of Trento, Italy. Currently, he is a post-doctoral research associate at Caltech. His research interests are structured representation learning.
[Uncaptioned image] T. Anderson Keller is a postdoctoral research fellow at the Kempner Institute at Harvard University. He completed his doctorate under the supervision of Max Welling at the University of Amsterdam. His current research focuses on structured representation learning, probabilistic generative modeling, and biologically plausible learning. His research explores ways to develop deep probabilistic generative models that are meaningfully structured with respect to observed, real-world transformations.
[Uncaptioned image] Yisong Yue is a professor of Computing and Mathematical Sciences at Caltech. My research interests lie primarily in machine learning, and span the entire theory-to-application spectrum from foundational advances all the way to deployment in real systems. I work closely with domain experts to understand the frontier challenges in applied machine learning, distill those challenges into mathematically precise formulations, and develop novel methods to tackle them.
[Uncaptioned image] Pietro Perona received the graduate degree in electrical engineering from the Università di Padova in 1985 and the PhD degree in electrical engineering and computer science from the University of California at Berkeley in 1990. After a postdoctoral fellowship at MIT in 1990-1991 he joined the faculty of the California Institute of Technology, Caltech in 1991, where he is now an Allen E. Puckett professor of electrical engineering and computation and neural systems. His current interests include visual recognition, modeling vision in biological systems, modeling and measuring behavior, and Visipedia. He has worked on anisotropic diffusion, multiresolution-multiorientation filtering, human texture perception and segmentation, dynamic vision, grouping, analysis of human motion, recognition of object categories, and modeling visual search.
[Uncaptioned image] Max Welling is a research chair in Machine Learning at the University of Amsterdam and a Distinguished Scientist at MSR. He is a fellow at the Canadian Institute for Advanced Research (CIFAR) and the European Lab for Learning and Intelligent Systems (ELLIS) where he also serves on the founding board. His previous appointments include VP at Qualcomm Technologies, professor at UC Irvine, postdoc at U. Toronto, and UCL under the supervision of Prof. Geoffrey Hinton, and postdoc at Caltech under the supervision of Prof. Pietro Perona. He finished his PhD in theoretical high energy physics under the supervision of Nobel laureate Prof. Gerard ‘t Hooft.